OpenStudy (anonymous):

simplify (1/x)/(2x+17) please help!

OpenStudy (mathteacher1729):

Hint: recognize that 2x+ 17 is really (2x+17) over 1. \[\huge \dfrac{\left(\dfrac{1}{x}\right)}{\left(\dfrac{2x+17}{1}\right)}\] Then "invert and multiply" the fraction in the denom (2x + 17) over 1. Hope this helps. :)